A certain scale for weighing food registers only weights above 6 pounds . a person who wanted to know the weight of 1 package each of chicken , beef, and turkey weighed every possible pair of these packages and got the following results.
The chicken and the beef weighed 7 pounds
The chicken and the turkey weighed 8 pounds
The beef and the turkey weighed 9 pounds
What is the weight of the package of turkey?
A 2 pounds b 3 pounds c 4 pounds d 5 pounds e 6 pounds

Answer by richwmiller(9143)

c+b=7,
c+t=8,
b+t=9
c=8-t
b=9-t
sub in the first equation
8-t+9-t=7
17-7=2t
10=2t
5=t
b=4 c=3